OpenStudy (anonymous):

Which answer represents the domain of the logarithmic function given below? F(x) = log9 x

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle {\rm F}(x)= {\rm Log}_9(x)}\)

OpenStudy (solomonzelman):

you can take a log of any positive number but \(\large\color{slate}{\displaystyle {\rm Log}_9(0)}\), or \(\large\color{slate}{\displaystyle {\rm Log}_9({\rm negative~~ number})}\) is undefined.

OpenStudy (solomonzelman):

the x in the log must be positive